math

You are asked to evaluate the following two projects for Boring Corporation. Use a discount rate of 10 percent. Use Appendix B.

 

Project X (DVDs
of the Weather Reports)
($14,000 Investment)


Project Y (Slow-Motion
Replays of Commercials)
($34,000 Investment)

Year   Cash Flow
YearCash Flow
1
$7,000

1
$17,000
2
5,000

2
10,000
3
6,000

3
11,000
4
5,600

4
13,000

 

a. Calculate the profitability index for project X. (Round "PV Factor" to 3 decimal places. Round the final answer to 2 decimal places.)

 

PI            Not attempted

 

b. Calculate the profitability index for project Y. (Round "PV Factor" to 3 decimal places. Round the final answer to 2 decimal places.)

 

PI             1.20 Correct

 

c. Using the NPV method combined with the PI approach, which project would you select? Use a discount rate of 10 percent.

 

  • Project Y
Answer #1

a)

Project X:

Profitability index = Present value / initial investment

Present value = 7,000 / (1 + 0.1)1 + 5,000 / (1 + 0.1)2 + 6,000 / (1 + 0.1)3 + 5,000 / (1 + 0.1)4

Present value = $18,418.8229

Profitability index = 18,418.8229 / 14,000

Profitability index = 1.32

b)

Project Y:

Profitability index = Present value / initial investment

Present value = 17,000 / (1 + 0.1)1 + 10,000 / (1 + 0.1)2 + 11,000 / (1 + 0.1)3 + 13,000 / (1 + 0.1)4

Present value = $40,862.6460

Profitability index = 40,862.6460 / 34,000

Profitability index = 1.20

c)

Project X as it has the higher profitability index